OpenStudy (anonymous):

Find the general solution of sin x + sin 2x + sin 3x + sin 4x = 0 ?

sam (.sam.):

Have you tried using identities to solve it?

sam (.sam.):

Like sin(2x)=2sin(x)cos(x)

OpenStudy (anonymous):

no

sam (.sam.):

For sin 3x , split it then use identity =sin(3x) =sin(2x+x) Use sin(A+B)=sin(A)cos(B)+cos(A)sin(B)

OpenStudy (anonymous):

sinx+sin2x+sin2xcosx+cos2xsinx+sin4x

OpenStudy (anonymous):

i have put the value in the given question

sam (.sam.):

as for sin4x as well, split it sin(2x+2x)

sam (.sam.):

then use sin(2x)=2sin(x)cos(x) for all

OpenStudy (anonymous):

so the given equation will becomes sinx+2sinxcosx+sin2xcosx+cos2xsinx+sin2xcos2x+cos2xsin2x

sam (.sam.):

yes then use sin(2x)=2sin(x)cos(x) for those with sin(2x)

sam (.sam.):

You should get something like this sin(x) + 2sin(x)cos(x) + [2sin(x)cos^2(x)+sin(x)cos(2x)] + 4sin(x)cos(x)cos(2x) = 0

sam (.sam.):

Then factor sin(x) out and you will get a solution

OpenStudy (anonymous):

sinx+2sinxcosx+2sinxcosxcosx+cos2xsinx+2sinxcosxcos2x+2cos2xsinxcosx=0

sam (.sam.):

Try simplify it you should get something like this sin(x) + 2sin(x)cos(x) + [2sin(x)cos^2(x)+sin(x)cos(2x)] + 4sin(x)cos(x)cos(2x) = 0

sam (.sam.):

Once you've done it then factor sin(x), sin(x)[1 + 2cos(x) + 2cos^2(x) + cos(2x) + 4cos(x)cos(2x)] = 0 You can say from here, sin(x)=0 then x = 2nπ for any integer n

OpenStudy (anonymous):

so x= 2nπ is the final answer